The difference of these two positive numbers is 9. And when you add 3 in each number the product will be 120. What is the answer using quadratic equation and factoring?
Answer by jorel1380(3719)

Let the second number be n. Then the first number is n-9. So:
(n+3)(n-6)=120
nē-3n-138=0
Using the quadratic formula, we get a positive value of 13.3427192823 for n. So, n-9=4.3427192823; which gives us our two numbers. ☺☺☺☺